OpenStudy (anonymous):

solve. 10=c/3-4+c/6 they gave me the answer and c=28 but they are asking me to explain that answer and I don't know how

OpenStudy (anonymous):

c=28

OpenStudy (anonymous):

Move the 4 to the other side of the equal sign 14=c/3+c/6 Change c/3 to 2C/6 (You need both denominators the same so you can add them) 14=2c/6+c/6 You can now add both together 14=3c/6 14=c/2 Multiply both sides by 2 28=c

OpenStudy (anonymous):

thank you